OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 610! If four times the number of one-step equations minus twice the number of two-step equations is equal to 340, how many one-step equations auditioned today? 260 340 350 670

OpenStudy (abb0t):

call the unknowns "x" and "y" so if the total number of one and two-step equations was 610 then: x + y = 610

OpenStudy (anonymous):

A.

OpenStudy (ddcamp):

let a = number of 1 step equations and b = number of 2 step equations a + b = 610 4a - 2b = 340 solve for a

OpenStudy (abb0t):

let "x" = onestep. so if its 4x the number of one-step, then that's 4x and twice the two-step equations would be 2y. 4x-2y = 340 and so forth.

OpenStudy (abb0t):

now you have a system. can you solve for them now?

OpenStudy (anonymous):

so it is A?

OpenStudy (abb0t):

for "a" or "x" ? you should be able to see the answer :)

OpenStudy (anonymous):

YES! A is the answer!!

OpenStudy (abb0t):

lol

OpenStudy (anonymous):

thank you!

OpenStudy (abb0t):

Guess you didn't learn anything, coleman :(